### v1

@@ Abstract
This demo paper reports on a new workflow GhostWriter that combines the use of Large Language Models and Knowledge Graphs (semantic artifacts) to support navigation through collections. Situated in the research area of Retrieval Augmented Generation, this specific workflow details the creation of local and adaptable chatbots. Based on the tool-suite EverythingData at the backend, GhostWriter provides an interface that enables querying and “chatting” with a collection. Applied iteratively, the workflow supports the information needs of researchers when interacting with a collection of papers, whether it be to gain an overview, to learn more about a specific concept and its context, and helps the researcher ultimately to refine their research question in a controlled way. We demonstrate the workflow for a collection of articles from the method data analysis journal published by GESIS – Leibniz-Institute for the Social Sciences. We also point to further application areas.
@@ 1 Introduction
Who would not love to have the possibility to navigate through a collection of documents by asking all kinds of questions to it? Traditionally, knowledge is communicated by language in a more or less codified way. On the less codified, natural language side, we find the development and application of Generative AI, which are primarily based on Large Language Models (LLM’s) and mimic chatting with an “expert" [1]. At the other end of the spectrum of how knowledge is communicated, we find Knowledge Organization Systems (KOS) [2] (and more specifically, semantic artifacts [3]). To be queried, these require some prior understanding of the structure in which the knowledge is codified. Think here about how one would have used a systematic catalog in a library in which books are indexed and ordered according to classification systems. Metaphorically, one could describe those two ends of the spectrum of how to engage with knowledge as “asking an expert" versus “asking a librarian".
@@
The motivation for this explorative research is multi-fold. First, it starts with being aware of the ambiguity of natural language. There is power in this ambiguity: It serves serendipity and enables the emergence of associations across contexts. But this ambiguity might also lead to possible misunderstandings. So, to compensate for the ambiguity of natural language, one might look in the direction of more defined, codified terminology as expressed in KOS. But terms from KOS also carry a burden. By their nature, they are defined in specific contexts and usually do not easily travel across knowledge domains. The experimental workflow and demo we present here aspire to combine the best of both ways to communicate knowledge when it comes to information retrieval. Scientifically, it is located in the area of Retrieval Augmented Generation (see 2). Another motivation for developing this workflow lies in the fact that many generative AI solutions are provided by large tech companies. They are trained on specific materials, the origin of which is not always known. Additionally, these new generations of AI are bases on stochastic methods, and so the results cannot be reproduced in a deterministic way. So-called ’hallucinations’, assertions that misrepresent facts and context, are one known effect of those stochastic models [4]. With this demo paper, we don’t totally abandon LLM’s and industry provided solutions, but we discuss how they can become embedded in workflows which can be controlled locally, so that AI is applied in a way following Open Science principles, and being reproducible and ethical.
@@
Our paper starts with a short overview of some related works in section 2. We introduce the use case in section 3. The description of the GhostWriter workflow and some demonstrations of its abilities form the core of this paper in section 4. The paper closes with some outlook in section 5.
@@ 2 Related work
Information Retrieval (IR) is the task to find specific information from a large collection of documents [5] given a query. This is not only an important step in the scientific endeavour to answer a specific research question, but also more generally a crucial way to reduce information and cognitive overload [6]. With the help of LLMs, Retrieval Augmented Generation (RAG) [7] solves this task in a modern approach, in which the documents, in particular the paragraphs, are stored in the vector database, together with their semantic embeddings. Once a query is proposed, it will be converted into the embedding space and matched with the embedding vectors of the paragraphs stored in the vector database by calculating the cosine similarity. The matched paragraphs are consequently summarized by the LLM, and the query is then answered with natural language.
With the RAG method, a query could be answered appropriately with less likelihood of triggering the hallucination of the LLMs, as long as the paragraphs can be matched precisely. However, for those questions that require a comprehensive understanding of the documents globally, especially in cases when the concepts are vital to answer a query, but they are located far apart in the documents, the standard RAG performs less well. As an improvement, the GraphRAG [8] targets the same kind of query and constructs the paragraphs in a graph structure, where the entities and relationships are extracted and are used as the vertices and edges of a graph, together with a text description generated by the LLM. This graph is then partitioned hierarchically into subgraphs, called the graph communities. To answer a query, the graph communities are summarized by the LLM. The LLM then scores whether this summarization answers the query. To generate the final answer, the summaries are ranked by the LLM-based score and concatenated until the token limit is reached. The GraphRAG performs well on the tasks that require the information located in different parts of documents. However, the graph might be sparse when the entities and relationships are not explicitly explained or paraphrased, and the connections will break. In contrast, our workflow is based on the RAG paradigm, but by using the ontology and knowledge graphs, we emphasize the generality between the entities and relationships to improve the matching accuracy in a wider range within the documents.
@@ 4.1 System Design
For this demo paper, we decided to present elements of the workflow schematically in Figure 1. In the upper part, we find a description of EverythingData part of the workflow, while the lower part represents the GhostWriter interface. EverythingData is a generic way to extract and enrich information from documents of any kind. After documents are split and paragraphs are extracted (1, in Figure 1), terms from these paragraphs are extracted (2) and those terms are further enriched (3). For each of these steps, embeddings are created, using a vector database in which those operations take place. As a result, we find terms which are already selected, weighted, extended according to the way the embeddings are executed, with the result that terms become contextualized further. In the lower part of Figure 1, we schematically summarize the GhostWriter interface. It starts with a query (already taken from our use case) for which information is retrieved (4). Eventually, the human-computer interface presents a natural language answer and pointers to possible source documents (5). This last step can be repeated, thus enabling an iterative way of further specifying the original query (chatting with the local collection). The data enrichment in EverythingData is enabled by the LLM-based term extraction and the knowledge graph & ontology-based term enrichment, while the HCI interface is achieved by the LLMs, which enables natural language. One interesting feature of this approach is that multilinguality is naturally part of it, due to the coupling to knowledge graphs which connect various language spaces (such as wikidata).
@@ 4.2 A detailed chat with method-data-analysis papers
To showcase our workflow, we created a collection with 100 articles from the mda journal [^6] Ghostwriter on MDA papers, betaversion, see https://gesis.now.museum, select collect collection mda. We then proposed questions to the interface and inspected the results. For this demo paper we choose to display the question: “explain male breadwinner model to me". The first response is shown in a snapshot from the website (see Figure 2). One can see that the question is explained in detail, and related articles are listed as well, together with a score of confidence. This indicates that this workflow can answer questions in an academic language and support further investigation into the collection.
@@
Figure 3: Finetuning the question on the “male breadwinner model"
@@ 5 Conclusion and Outlook
This demo paper bases on experiences gained in a series of project-based explorations concerning information processes in the social sciences and humanities, and more specifically in the area cultural heritage and their information infrastructural needs, such as the European project MuseIT [^7] https://www.muse-it.eu [18] and the Dutch research infrastructure project ODISSEI [^8] https://odissei-data.nl. As we demonstrated in Section 4, the workflow belongs to the RAG approach and relies on entity extraction, where terms are annotated with semantic meaning by mapping them into knowledge graphs build from controlled vocabularies. These ways to connect vector spaces (LLM’s) to a knowledge graph (structured knowledge, KOS) open possibilities to enhance the semantic meaning for information retrieval tools. Linking entities to even more knowledge graph representations (such as wikidata) brings the possibility of multilingual search as one immediate benefit of such enrichment. As demonstrated in the case of chatting with (some) mda papers, any query input starts a kind of distant reading machinery which returns with an answer (using natural language) together with pointers to the sources from the original collection supporting this provisional answer. This way, it mimics a close-reading level and still enables the user to continue her/his close-reading journey. The GhostWriter workflow works for any “collection". You can create a “semantic index" and a local chatbot for collections from webpages/websites, RSS feeds, but also from trusted data repositories such as Dataverse instances. (see DataChat [^9] https://github.com/gdcc/datachat). In essence, you enable local chatting with papers – this way pointing a human to specific texts which one can then engage with by way of close-reading. You confine the ‘search space’ by placing your question (and the collection) into a particular area in the networked space of scientific knowledge. We also demonstrated that you gain information beyond what is explicitly expressed in the text and annotated in the metadata (by creating associations on the level of natural language and on the level of Knowledge Organisation Systems). It is like you start chatting with the experts or invisible colleges behind the papers. But, there are also many preconditions to make such a workflow work. It starts with API’s from which collections can be harvested, and API’s which provide KOS services (for the risks of KOS services see e.g. [19]). Well-curated metadata and digitized full text are another precondition. Implementation of the Croissant Standard [15] as a new standard for datasets for machine-learning is another precondition.
@@
So in short, while this technology offers great opportunities, it is far from being a ’ready-from-the-shelf-implementable’ software. Having said this, one large application field is collections of documents from the Cultural Heritage - for instance, the many collections that Galleries, Libraries, Archives and Musea hold [20] contain. Developing skills to develop your workflows and to tailor them to your own needs is a very important way to empower Global Open Research Commons [21], a way to execute Open Science, even more needed today.

### v2

@@ Abstract
This demo paper reports on a new workflow GhostWriter that combines the use of Large Language Models and Knowledge Graphs (semantic artifacts) to support navigation through collections. Situated in the research area of Retrieval Augmented Generation, this specific workflow represents the creation of local and adaptable chatbots. Based on the tool-suite EverythingData at the backend, GhostWriter provides an interface that enables querying and “chatting” with a collection. Applied iteratively, the workflow supports the information needs of researchers when interacting with a collection of papers, whether it be to gain an overview, to learn more about a specific concept and its context, and helps the researcher ultimately to refine their research question in a controlled way. We demonstrate t he workflow for a collection of articles from the method data analysis journal published by GESIS – Leibniz-Institute for the Social Sciences. We also point to further application areas.
@@ 1 Introduction
Who would not love to have the possibility to navigate through a collection of documents by asking all kinds of questions to it? Traditionally, knowledge is communicated by language in a more or less codified way. On the less codified, natural language side, we find the development and application of Generative AI, which are primarily based on Large Language Models (LLMs) and mimic chatting with an “expert" [1]. At the other end of the spectrum of how knowledge is communicated, we find Knowledge Organization Systems (KOS) [2] (and more specifically, semantic artifacts [3]). To be queried, the latter require some prior understanding of the structure in which the knowledge is codified. Think here about how one would have used a systematic catalog in a library in which books are indexed and ordered according to classification systems. Metaphorically, one could describe those two ends of the spectrum of how to engage with knowledge as “asking an expert" versus “asking a librarian".
@@
The motivation for this explorative research is multifold. First, it starts with being aware of the ambiguity of natural language. There is power in this ambiguity: It serves serendipity and enables the emergence of associations across contexts. But this ambiguity might also lead to possible misunderstandings. So, to compensate for the ambiguity of natural language, one might look in the direction of more defined, codified terminology as expressed in KOS. But terms from KOS also carry a burden. By their nature, they are defined in specific contexts and usually do not easily travel across knowledge domains. The experimental workflow and demo we present here aspire to combine the best of both ways to communicate knowledge when it comes to information retrieval. This demo paper is scientifically located in the area of Retrieval Augmented Generation (see 2). An additional motivation for developing this workflow lies in the fact that many generative AI solutions are provided by large tech companies. They are trained on specific materials, the origin of which is not always known. Additionally, these new generations of AI are based on stochastic methods, and so the results cannot be reproduced in a deterministic way. The so-called ’hallucinations’, assertions that misrepresent facts and context, are one known effect of these stochastic models [4]. Despite these features, the power of Large Laguage Models (LLMs) is increasingly unfolded in all kind of research. The goal of the development of a fully functional prototype as reported in this demo paper was also to explore how data streams can be effectively connected to LLMs on a smaller scale, ensuring that relevant data segments are delivered to the model at the right time. Knowledge graphs appear to be a promising approach for sharing structured information across different models, as they include provenance and data origin, which are essential for trust and context, and combined with LLM’s can enhance information retrieval. With this demo paper, we discuss how LLM’s and industry provided solutions can become embedded in workflows which can be controlled locally, so that AI is applied in a way following Open Science principles, and being reproducible and ethical.
@@
Our paper starts with a short overview of some related works in section 2. We introduce the use case in section 3. The description of the GhostWriter workflow and some demonstrations of its abilities form the core of this paper in section 4. The paper closes with conclusions and an outlook into future research topics in section 5.
@@ 2 Related work
Information Retrieval (IR) is the task of finding specific information from a large collection of documents [5] given a query. This is not only an important step in the scientific endeavor to answer a specific research question, but also, more generally, a crucial way to reduce information overload and cognitive overload [6]. With the help of LLMs, augmented search generation (RAG) [7] solves this task in a modern approach, in which documents, and particular paragraphs from them, are stored in the vector database, together with their semantic embeddings. Once a query is proposed, it will be converted into the embedding space and matched with the embedding vectors of the paragraphs stored in the vector database by calculating the cosine similarity. The matched paragraphs are consequently summarized by the LLM, and the query is then answered with natural language.
With the RAG method, a query could be answered appropriately with less likelihood of triggering the hallucination of the LLMs, as long as the paragraphs can be matched precisely. But there are situations in which the standard RAG performs less well. For instance, when questions require a comprehensive understanding of the documents on a more global scale, and especially in cases when the concepts which are vital to answer a query are located far apart in the documents. As an improvement, the GraphRAG [8] targets the same kind of query and constructs the paragraphs in a graph structure, where the entities and relationships are extracted and are used as the vertices and edges of a graph, together with a text description generated by the LLM. This graph is then partitioned hierarchically into subgraphs, called the graph communities. To answer a query, the graph communities are summarized by the LLM. The LLM toolbox contains scores which indicate whether this summarization answers the query. To generate the final answer, the summaries (of the graph communities) are ranked to the LLM-based score and concatenated until the token limit is reached. The GraphRAG method has been proven to performs well on tasks that require the information to be located in different parts of documents. However, the graph might be sparse when the entities and relationships are not explicitly explained or paraphrased, and hence the connections will break. In contrast, the workflow of this demo paper while based on the RAG paradigm, uses ontology and knowledge graphs to eventually define graph communities. This way, the workflow emphasizes the generality between entities and relationships in order to improve the matching accuracy in a wider range within the documents.
@@ 4.1 System Design
For this demo paper, we decided to present elements of the workflow schematically in Figure 1. In the upper part, we find a description of EverythingData [^6] https://github.com/Dans-labs/everythingdata part of the workflow, while the lower part represents the GhostWriter interface. EverythingData is a generic way to extract and enrich information from documents of any kind. After documents are split and paragraphs are extracted (1, in Figure 1), terms from these paragraphs are extracted (2) and those terms are further enriched (3). For each of these steps, embeddings are created, using a vector database in which those operations take place. As a result, we find terms which are already selected, weighted, extended according to the way the embeddings are executed, with the result that terms become contextualized further. In the lower part of Figure 1, we schematically summarize the GhostWriter interface. It starts with a query (already taken from our use case) for which information is retrieved (4). Eventually, the human-computer interface presents a natural language answer and pointers to possible source documents (5). This last step can be repeated, thus enabling an iterative way of further specifying the original query (chatting with the local collection). The data enrichment in EverythingData is enabled by the LLM-based term extraction and the knowledge graph & ontology-based term enrichment, while the HCI interface is achieved by the LLMs, which enables natural language. One interesting feature of this approach is that multilinguality is naturally part of it, due to the coupling to knowledge graphs which connect various language spaces (such as wikidata).
@@ 4.2 A detailed chat with method-data-analysis papers
To showcase our workflow, we created a collection with 100 articles from the mda journal [^7] Ghostwriter on MDA papers, betaversion, see https://gesis.now.museum, select collect collection mda. We then proposed questions to the interface and inspected the results. For this demo paper, we choose to display the question: “explain male breadwinner model to me". The first response is shown in a snapshot from the website (see Figure 2). One can see that the question is explained in detail, and related articles are listed as well, together with a score of confidence (calculated inside the workflow). This indicates that this workflow can answer questions in an academic language and support further investigation into the collection.
@@
Figure 3: Finetuning the question on the “male breadwinner model"
@@ 5 Conclusion and Outlook
This demo paper bases on experiences gained in a series of project-based explorations concerning information processes in the social sciences and humanities, and more specifically in the area cultural heritage and their information infrastructural needs, such as the European project MuseIT [^8] https://www.muse-it.eu [18] and the Dutch research infrastructure project ODISSEI [^9] https://odissei-data.nl.
As we demonstrated in Section 4, the workflow belongs to the RAG approach and relies on entity extraction, where terms are annotated with semantic meaning by mapping them into knowledge graphs built from controlled vocabularies. These ways to connect vector spaces (LLM’s) to a knowledge graph (structured knowledge, KOS) open possibilities to enhance the semantic meaning for information retrieval tools.
The challenge with this approach - which in essence is combining semantically related pieces of text - is achieving a coherent interpretation of the context. It often fails when contradictions or differing points of view are present in the source collection. Additionally, semantic similarity does not guarantee logical consistency, and models may incorrectly link unrelated content due to superficial lexical overlap. There’s also a risk of the wrong understanding of the context, where nuanced or domain-specific meanings are lost.
As a possible future solution, the current ranking algorithm should be revised to first cluster and rank document fragments at the document level. This means grouping semantically similar fragments by their source document before evaluating their relevance. By doing this, we avoid mixing contradictory content from different documents and preserve contextual coherence. Once ranked, selected clusters can be passed to the LLM, along with explicit metadata about provenance such as paper titles, authors, publication dates, and source repositories. This enhances the reliability of the generated responses by making the origins of information traceable and allowing the model to distinguish between perspectives from different sources. Linking entities to even more knowledge graph representations (such as wikidata) brings the possibility of multilingual search as one immediate benefit of such enrichment.
As demonstrated in the case of chatting with (some) mda papers, any query input starts a kind of distant reading machinery which returns with an answer (using natural language) together with pointers to the sources from the original collection supporting this provisional answer. This way, it mimics a close-reading level and still enables the user to continue her/his close-reading journey.
@@
The GhostWriter workflow works for any “collection". You can create a “semantic index" and a local chatbot for collections from webpages/websites, RSS feeds, but also from trusted data repositories such as Dataverse instances. (see DataChat [^10] https://github.com/gdcc/datachat). In essence, you enable local chatting with papers – this way pointing a human to specific texts which one can then engage with by way of close-reading. You confine the ‘search space’ by placing your question (and the collection) into a particular area in the networked space of scientific knowledge. We also demonstrated that you gain information beyond what is explicitly expressed in the text and annotated in the metadata (by creating associations on the level of natural language and on the level of Knowledge Organisation Systems). It is like you start chatting with the experts or invisible colleges behind the papers. But, there are also many preconditions to make such a workflow work. It starts with API’s from which collections can be harvested, and API’s which provide KOS services (for the risks of KOS services see e.g. [19]). Well-curated metadata and digitized full text are another precondition. Implementation of the Croissant Standard [15] as a new standard for datasets for machine-learning is another precondition. As part of our future work, we have developed the Model Context Protocol (MCP) [^11] https://mcp.dataverse.org and initiated the practice of depositing metadata for papers separately in the Dataverse data repository. This setup allows MCP to be used as an automated factual query source. The approach appears more reliable than other approaches, as it retrieves up-to-date metadata directly from data sources and enables verification of the information provided. It’s also currently being deployed as “Ask Dataverse” services and tested by Harvard IQSS on https://ask.dataverse.org. So in short, while this technology offers great opportunities, it is far from being a ’ready-from-the-shelf-implementable’ software. It is still very much in development.
@@
Having said this, one large application field is collections of documents from the Cultural Heritage - for instance, the many collections that Galleries, Libraries, Archives and Musea hold [20] contain. Developing skills to develop your workflows and to tailor them to your own needs is a very important way to empower Global Open Research Commons [21], a way to execute Open Science, even more needed today.